OpenStudy (anonymous):

solve -7(11-4w)-15w=-27-(50-13w)

OpenStudy (anonymous):

\[-77 + 28w - 15w = -27-50+13w\] \[-77+13w=-77+13w\] They are both the same equation, so you can't find the solution.

OpenStudy (anonymous):

ok thanks